# Reconstructing the recombination history by combining early and late cosmological probes

### v2

@@ I Introduction
Responses of CMB angular power spectra to changes in recombination have previously been exploited to constrain elements of the recombination process. For example, the Planck collaboration used measurements of the temperature and polarization power spectra to constrain the atomic $2s\to 1s$ two-photon transition with 7% uncertainty [6]. Additionally, phenomenological modifications to the recombination history have been constrained under the assumption that the angular power spectra respond linearly to changes in $X_{\rm e}(z)$, what we will refer to henceforth as a linear response approximation, or LRA [7, 1, 8]. This sensitivity has also been used to constrain physical models with energy injection into the primordial plasma [9, 10, 11, 12], varying fundamental constants [13, 14, 15, 16], and small-scale clumping of baryons during recombination [17, 18, 19, 20]. The CMB is sensitive to all of these models primarily through the changes induced in the recombination process around $z\simeq 10^{3}$.
@@
Whether or not we use an LRA impacts the inferred values and uncertainties of cosmological parameters. PCA-based approaches to perturbatively modified recombination, which make this approximation, generally find little to no shift in parameter values and negligible loosening of constraints, even when up to four eigenmodes are retained [8, 1]. However, it should be remarked that the PCA has not been applied directly to CMB+SN data combinations, which is expected to hint towards modified a recombination history.
@@ II.1 Details of the modified recombination model
II.2 Accelerated inference using an emulator
@@
[…]
The emulator had 6 hidden layers with 512 fully connected nodes per layer. We trained the emulator for 150 epochs with a batch size of 256 using a mean square error loss function. We terminated the iterative process once the 95th percentile errors on angular power spectra were below 1%. As an additional post hoc assessment of the emulator accuracy, we importance sampled an MCMC chain produced with this emulator, with weights set by the likelihood as computed using CLASS-calculated power spectra. This re-weighted distribution showed little change as compared to the original distribution obtained using the emulator, indicating that the emulator is sufficiently accurate for inference in regions of appreciable likelihood. A comparison of the original and re-weighted distributions in the $H_{0}-\omega_{b}$ plane is shown in Fig. 3. We also show the 95th percentile $\Delta\chi^{2}$ values for points drawn from the posterior in this region. See Appendix A for further details on the emulator and additional error quantification.
@@ III.1.3 Diffusion damping
For more on the performance of this approximation through decoupling, see [54]. Of course in Einstein-Boltzmann solvers the tight coupling approximation is abandoned sufficiently early to preserve the accuracy of model power spectra [e.g. 33].
@@ III.2 Reconstruction from Planck CMB data
The increased uncertainty in standard parameters is in contrast to previous perturbative phenomenological studies which have found that constraints are not greatly loosened [1, 7]. The increased uncertainty is a result of our choice to avoid the LRA and to instead fully treat the impact of a modified $X_{\rm e}(z)$ on the final radiation anisotropies. Our results are consistent with other studies focused on particular physical models which do see an effect on parameter constraints [e.g., 15], albeit in a model-independent way.
III.3 Constraints from Planck and BAO data
The same initial conditions and dynamics that lead to a series of peaks in the CMB power spectrum result in a series of peaks in the matter power spectrum, and, equivalently, a peak in the matter correlation function [56, 57, 58, 59, 47]. This BAO feature is observed as an excess of clustering in the large-scale galaxy distribution on comoving scales roughly equal to the sound horizon at the redshift $z_{\rm drag}$ at which baryons decouple from the radiation. This occurs at a slightly later redshift than last scattering due to the low baryon-to-photon ratio, and is precisely defined conventionally as the redshift at which the baryon optical depth passes unity. The sound horizon at this epoch is
@@ III.3 Constraints from Planck and BAO data
We use measurements of the BAO feature from the eBOSS DR16 release [47], including luminous red galaxies (LRG), emission line galaxies (ELG), quasars, and Lyman-$\alpha$ forest samples. We also include the BOSS galaxy sample [60] as well as the SDSS main galaxy sample (MGS). In total these data provide measurements of the low redshift distances in Eq. 14 in seven bins between $2.330.15<z<$ 2.33. We also include redshift space distortion measurements in our analysis. Modern BAO measurements use pre-computed templates of the matter power spectrum exhibiting the BAO feature and using an assumed cosmological model, which are then re-scaled to fit the observed power spectrum. Following the analysis of Bernal et al. [61], we verified that the ModRec model does not introduce additional contributions which may complicate the template-based determination.
@@
In Fig. 8, we show these BAO measurements along with low-redshift distances computed using models from our Planck-only chain. For a given curve, we bin models drawn from the chain centered on integer values of $H_{0}$ with unit bin width, in units of km/s/Mpc, and plot the mean of the resulting set of curves. Each curve is therefore representative of models that have $H_{0}$ near the labeled values. We only show curves representing $H_{0}$ values within the 95% confidence limits for $H_{0}$ in the Planck chain.
@@
The results of the Planck+BAO MCMC chain are shown in the red contours in Fig. 6. Consistent with the discussion above, constraints on $H_{0}$ are tightened from $H_{0}=69.2\pm 6.6$ in the case of Planck data alone to $H_{0}=68.9\pm 1.1$ when including BAO data. This is consistent with the conclusions in [15] for a varying $m_{\rm e}$ model.
@@ IV The $H_{0}$ tension with modified recombination
Using our reconstructions of the ionization history, we examine the extent to which non-perturbatively modified recombination can accommodate higher values of $H_{0}$. This extends existing phenomenological studies [e.g., 21, 8] to non-perturbative changes to the recombination history in a physical model-independent way. We are also able to study where in redshift space modifications to $X_{\rm e}$ become important in terms of resolving the tension and identify specific modifications which do a good job of alleviating the tension while maintaining a fit to the data. This may be useful to model builders seeking to alleviate the $H_{0}$ tension through models which primarily affect the ionization history. We begin with a brief review of why modified recombination is able to alleviate the tension at all, and then study which recombination histories can achieve this while maintaining a good fit to the data.
@@ IV.2 High $H_{0}$ recombination histories
We now examine in more detail those recombination histories which can alleviate the Hubble tension. As a criterion to identify models which maintain a fit to the data, we select models from the chains with $\Delta\chi^{2}\equiv\chi^{2}_{{\tt ModRec}}-\chi^{2}_{\Lambda{\rm CDM,bf}}<0$. Here, $\chi^{2}_{\Lambda{\rm CDM,bf}}$ is $\chi^{2}$ of the best fit to the Planck+BAO data combination, assuming $\Lambda$CDM.
@@
In Fig. 10, we show the ionization fractions for this selection of models. Allowing for non-perturbative changes to recombination, a wide range of $H_{0}$ values are consistent with Planck data, including values as high as the SH0ES [62] measurement. For these high $H_{0}$ models, the required departures from the standard recombination scenario are as large as $\sim 40\%$ near $z\sim 900$, and have an oscillatory redshift dependence. We note that a similar feature was identified by Lee et al. [22] in the context of perturbative changes introduced by a time-varying $m_{\rm e}$.
@@
As expected following the discussion in Sec. III, including BAO data restricts the range of $H_{0}$ values which can be obtained while maintaining a good fit to all of the data. The estimate for $H_{0}$ from this chain is $H_{0}=68.9\pm 1.1$ km/s/Mpc, indicating a reduction of the tension to a $2.7\sigma$ level. The ionization fraction is restricted to near its standard value, with the maximum departure among models satisfying our selection criterion being around 15% near $z=900$. This restriction is once again a consequence of the sensitivity of BAO data to low-redshift distance measures, as highlighted in Fig. 8 and Sec. III [see also the discussion in 15].
@@ IV.4 Inclusion of SH0ES data
The final column of Fig. 5 shows our reconstruction of recombination quantities when SH0ES data are included. These recombination histories differ substantially from the $\Lambda$CDM predictions, with the means for $\tilde{q}_{3}$,$\tilde{q}_{4}$, and $\tilde{q}_{5}$ being $3.3\sigma$, $3.1\sigma$, and $3.3\sigma$ away from 0 deviation respectively. One interpretation of this is that within our model space, if distance ladder based determinations of $H_{0}$ are taken at face value, then Planck and BAO data favor a non-standard recombination over the redshift range $933\lesssim z\lesssim 1200$. This is an analogous result to previous studies which have also identified that non-standard recombination-era physics is preferred, if one includes the SH0ES result in the analysis [e.g. 15, 75, 22]. Our result shows this independently of the assumed alternative model, and also identifies the recombination history that best accommodates both early and late universe measurements.
@@
The mean visibility function from our Planck+BAO+SH0ES chain peaks at a slightly higher redshift relative to the $\Lambda$CDM visibility function (computed using the Planck mean parameters), and is also skewed towards higher redshifts. As examples of physical models which recover similar features, in Fig. 12, we show the mean $g(z)$ from the SH0ES chain, the $\Lambda$CDM $g(z)$, as well as $g(z)$ arising from a model with primordial magnetic fields and varying electron mass. It is apparent that the latter exhibit a similar level of skewness and even kurtosis (see Table 3). However, for the PMF scenario, a number of approximations are made in the calculation of the related visibility function, which is based on computing the averaged recombination history including only second order effects of baryon-density perturbations on the recombination dynamics [e.g., 17, 18, for more details of the model]. Indeed the non-linear nature of the ionization history responses to baryon density fluctuations suggests that this may be omitting additional effects, which in a more rigorous treatment could manifest in enhanced higher order moments of the Thomson visibility function, an aspect that may be worth exploring in the future.
@@ V Forecasts for SPT-3G
Figure 14: Power spectra residuals between predictions from the ModRec model and the Planck mean cosmology, taken as the reference cosmology. Also shown are binned data from Planck, as well as the forecasted bandpower errors for SPT-3G in the Main-1500 configuration. Each have been binned with $\Delta\ell=75$. The power spectra correspond to models having the labeled values of $H_{0}$, in the same manner as Fig. 8.
Figure 15: Fractional uncertainties in our recombination reconstructions for the Planck, Planck+BAO, and Planck+BAO+SPT data combinations. Dashed lines indicated the $68\%$ and $95\%$ confidence intervals. SPT data will decrease the our reconstruction uncertainties by roughly a factor of 2, relatively uniformly across the redshift range in question. The fiducial cosmology for the SPT mock likelihood was taken to be the Planck mean cosmology.
@@
$\log{\mathcal{L}}=\left(\mathbf{C}_{\ell}-\mathbf{C}_{\ell}^{\rm fid}\right)^{\intercal}\Sigma_{\ell}\left(\mathbf{C}_{\ell}-\mathbf{C}_{\ell}^{\rm fid}\right)$ ) (16)
where $\mathbf{C}_{\ell}=\left(C_{\ell}^{TT},C_{\ell}^{EE},C_{\ell}^{TE},C_{\ell}^{\phi\phi}\right)$ and the Planck 2018 mean cosmology is taken as the fiducial. The covariance matrix $\Sigma_{\ell}$ incorporates the effects of both instrumental noise and foregrounds and is constructed in the manner described in [76]. We use noise levels corresponding to the Main-1500 configuration of SPT-3G, representing 5 years of observations on a sky area of 1500 $\rm deg^{2}$. This corresponds to the data that has already been collected since SPT-3G began observations in 2019. As the observation regions for Planck and SPT-3G are not independent, we remove large scale modes from the mock SPT-3G data to avoid double counting information. The $\ell_{\rm min}$ threshold was determined as roughly the multipole at which the signal-to-noise for SPT-3G is twice that of Planck. This approach was validated in Prabhu et al. [76], and was found to give results less than 8% worse than a full Fisher accounting for the overlapping sky regions. The multipole ranges we used were $\ell\in[2000,3500],[1000,3500],[750,4000],[30,4000]$ for TT, TE, EE, and $\phi\phi$ respectively.
@@
In Fig. 14, we show power spectra predictions from well fitting models which have a range of values for $H_{0}$; these are the same models presented as dashed lines in in Figs. 10 and  11. Also presented are current Planck data, as well as binned bandpower errors for SPT-3G, with bins of $\Delta\ell=50$ in each case. The data presented reflect the mock data used for these forecasts.
Even for $H_{0}\approx 68$ km/s/Mpc, we see that the ModRec model power spectrum predictions differ from the $\Lambda$CDM power spectrum predictions. The new freedom introduced in recombination is able to exploit some feature of the data to improve the fit relative to the case of standard recombination. This is also reflected by the fact that the median recombination history for the Planck chain in Fig. 5 differs from the standard history.
For these models, there are deviations from $\Lambda$CDM at the level of a few percent in the damping tail for the $TT$ and $EE$ spectra. Additionally, there is a qualitative change in the shape of the predicted power spectrum for models with $H_{0}\approx 68$ km/s/Mpc. In the TT spectrum, for models with values above this, the predictions between $2500\lesssim\ell\lesssim 3500$ collapse to roughly the same value. There is also a change in behavior for the EE spectrum, but it is different in nature.
The largest deviation from the $\Lambda$CDM reference prediction can be seen in the temperature and polarization cross-spectrum, which is consequently where SPT-3G data has the largest impact. This highlights the importance of precise measurements of the temperature and polarization power spectra at small scales for constraining alternative models of the recombination era. SPT-3G measurements will be sensitive enough to restrict the range of possible recombination histories, or to detect many such deviations should they be present.

### v3

@@ I Introduction
Responses of CMB angular power spectra to changes in recombination have previously been exploited to constrain elements of the recombination process. For example, the Planck collaboration used measurements of the temperature and polarization power spectra to constrain the atomic $2s\to 1s$ two-photon transition with 7% uncertainty [6]. Additionally, phenomenological modifications to the recombination history have been constrained under the assumption that the angular power spectra respond linearly to changes in $X_{\rm e}(z)$, what we will refer to henceforth as a linear response approximation, or LRA [7, 1, 8]. This sensitivity has also been used to constrain physical models with energy injection into the primordial plasma [9, 10, 11, 12, 13], varying fundamental constants [14, 15, 16, 17], and small-scale clumping of baryons during recombination [18, 19, 20, 21]. The CMB is sensitive to all of these models primarily through the changes induced in the recombination process around $z\simeq 10^{3}$.
@@
Whether or not we use an LRA impacts the inferred values and uncertainties of cosmological parameters. PCA-based approaches to perturbatively modified recombination, which make this approximation, generally find little to no shift in parameter values and negligible loosening of constraints, even when up to four eigenmodes are retained [8, 1]. However, it should be remarked that a PCA has not been applied directly to the combination of CMB and Cepheid-calibrated supernovae data, and doing so might hint towards a modified recombination history.
@@ II.2 Choice of control points
II.3 Accelerated inference using an emulator
@@
These quantities are concatenated to form the output vector which the neural network is trained to reproduce, given the corresponding input vector of the 13 model parameters, and this vector is normalized using the “standardization” option in CONNECT [see 24]. The neural network had 6 hidden layers with 512 fully connected nodes per layer. We trained for 150 epochs with a batch size of 256 using a mean square error loss function. We terminated the iterative process once the 95th percentile errors on angular power spectra were below 1%. As an additional post hoc assessment of the emulator accuracy, we importance sampled an MCMC chain produced with this emulator, with weights set by the likelihood as computed using CLASS-calculated power spectra. This re-weighted distribution showed little change as compared to the original distribution obtained using the emulator, indicating that the emulator is sufficiently accurate for inference in regions of appreciable likelihood. A comparison of the original and re-weighted distributions in the $H_{0}-\omega_{b}$ plane is shown in Fig. 3. We also show the 95th percentile $\Delta\chi^{2}$ values for points drawn from the posterior in this region. See Appendix A for further details on the emulator and additional error quantification.
The final training set contained 182640 models, of which 5% were reserved for test data. These are the only CLASS executions required for the entire analysis, which is the same order of magnitude as the number of accepted steps in a single chain within the ModRec model, although additional computations would be required for rejected steps. Furthermore, since we are emulating the theoretical quantities directly as opposed to the likelihoods, the same emulator can be used with a variety of data combinations. As long as some observable is included among or can be computed from the emulated quantities, the emulator can be used in combination with these data at no additional cost. Avoiding these computations with the trained emulator represents a significant acceleration of the inference process.
@@ III.1.3 Diffusion damping
For more on the performance of this approximation through decoupling, see [55]. Of course in Einstein-Boltzmann solvers the tight coupling approximation is abandoned sufficiently early to preserve the accuracy of model power spectra [e.g. 34].
@@ III.2 Reconstruction from Planck CMB data
The increased uncertainty in standard parameters is in contrast to previous perturbative phenomenological studies which have found that constraints are not greatly loosened [1, 7]. The increased uncertainty is a result of our choice to avoid the LRA and to instead fully treat the impact of a modified $X_{\rm e}(z)$ on the final radiation anisotropies. Our results are consistent with other studies focused on particular physical models which do see an effect on parameter constraints [e.g., 16], albeit in a model-independent way.
III.3 Reconstruction from CMB and BAO data
[…]
@@ III.3 Reconstruction from CMB and BAO data
We use measurements of the BAO feature from the eBOSS DR16 release [48], including luminous red galaxies (LRG), emission line galaxies (ELG), quasars, and Lyman-$\alpha$ forest samples. We also include the BOSS galaxy sample [61] as well as the SDSS main galaxy sample (MGS). In total these data provide measurements of the low redshift distances in Eq. 14 in seven bins between $2.330.15<z<$ 2.33. We include redshift space distortion measurements in our analysis, although we find that they are not very constraining. The robustness of these measurements in the presence of a modified recombination era is discussed in Appendix B.
The middle column of Fig. 5 shows the reconstructions of the ionization history using Planck+BAO data. The inclusion of BAO data sharpens the reconstructions and correspondingly tightens cosmological parameter constraints, which are shown in the red contours in Fig. 6. Consistent with the discussion above, constraints on $H_{0}$ are tightened from $H_{0}=69.2\pm 6.6$ km/s/Mpc in the case of Planck data alone to $H_{0}=68.9\pm 1.1$ km/s/Mpc when including BAO data. This is consistent with the conclusions in [16] for a varying $m_{\rm e}$ model.
@@
In Fig. 8, we show the BAO measurements along with low-redshift distances computed using models from our Planck-only chain. For a given curve, we bin models drawn from the chain centered on integer values of $H_{0}$ with unit bin width, in units of km/s/Mpc, and plot the mean of the resulting set of curves. Each curve is therefore representative of models that have $H_{0}$ near the labeled values. We only show curves representing $H_{0}$ values within the 95% confidence limits for $H_{0}$ in the Planck chain.
@@ IV The $H_{0}$ tension with modified recombination
Using our reconstructions of the ionization history, we examine the extent to which non-perturbatively modified recombination can accommodate higher values of $H_{0}$. This extends existing phenomenological studies [e.g., 22, 8] to non-perturbative changes to the recombination history in a physical model-independent way. We are also able to study where in redshift space modifications to $X_{\rm e}$ become important in terms of resolving the tension and identify specific modifications which do a good job of alleviating the tension while maintaining a fit to the data. This may be useful to model builders seeking to alleviate the $H_{0}$ tension through models which primarily affect the ionization history. We begin with a brief review of why modified recombination is able to alleviate the tension at all, and then study which recombination histories can achieve this while maintaining a good fit to the data.
@@ IV.2 High $H_{0}$ recombination histories
We now examine in more detail those recombination histories which can alleviate the Hubble tension. As a criterion to identify models which maintain a fit to the data, we select models from the chains with $\Delta\chi^{2}\equiv\chi^{2}_{\texttt{ModRec}}-\chi^{2}_{\Lambda{\rm CDM,bf}}<0$. Here, $\chi^{2}_{\Lambda{\rm CDM,bf}}$ is $\chi^{2}$ of the best fit to the Planck+BAO data combination, assuming $\Lambda$CDM.
@@
In Fig. 10, we show the ionization fractions for this selection of models. Allowing for non-perturbative changes to recombination, a wide range of $H_{0}$ values are consistent with Planck data, including values as high as the SH0ES [62] measurement. For these high $H_{0}$ models, the required departures from the standard recombination scenario are as large as $\sim 40\%$ near $z\sim 900$, and have an oscillatory redshift dependence. We note that a similar feature was identified by Lee et al. [23] in the context of perturbative changes introduced by a time-varying $m_{\rm e}$. However, these oscillations are not an essential feature to the overall fit: smoothing them in an appropriate manner degrades the fit by $\Delta\chi^{2}\lesssim 5$ [71].
@@
As expected following the discussion in Sec. III, including BAO data restricts the range of $H_{0}$ values which can be obtained while maintaining a good fit to all of the data. The estimate for $H_{0}$ from this chain is $H_{0}=68.9\pm 1.1$ km/s/Mpc, indicating a reduction of the tension to a $2.7\sigma$ level. The ionization fraction is restricted to near its standard value, with the maximum departure among models satisfying our selection criterion being around 15% near $z=900$. This restriction is once again a consequence of the sensitivity of BAO data to low-redshift distance measures, as highlighted in Fig. 8 and Sec. III [see also the discussion in 16].
@@ IV.4 Inclusion of SH0ES data
The final column of Fig. 5 shows our reconstruction of recombination quantities when SH0ES data are included. These recombination histories differ substantially from the $\Lambda$CDM predictions, with the means for $\tilde{q}_{3}$,$\tilde{q}_{4}$, and $\tilde{q}_{5}$ being $3.3\sigma$, $3.1\sigma$, and $3.3\sigma$ away from 0 deviation respectively. One interpretation of this is that within our model space, if distance ladder based determinations of $H_{0}$ are taken at face value, then Planck and BAO data favor a non-standard recombination over the redshift range $933\lesssim z\lesssim 1200$. This is an analogous result to previous studies which have also identified that non-standard recombination-era physics is preferred, if one includes the SH0ES result in the analysis [e.g. 16, 76, 23]. Our result shows this independently of the assumed alternative model, and also identifies the recombination history that best accommodates both early and late universe measurements.
@@
The mean visibility function from our Planck+BAO+SH0ES chain peaks at a slightly higher redshift relative to the $\Lambda$CDM visibility function (computed using the Planck mean parameters), and is also skewed towards higher redshifts. As examples of physical models which recover similar features, in Fig. 12, we show the mean $g(z)$ from the SH0ES chain, the $\Lambda$CDM $g(z)$, as well as $g(z)$ arising from a model with primordial magnetic fields and varying electron mass. It is apparent that the latter exhibit a similar level of skewness and even kurtosis (see Table 3). However, for the PMF scenario, a number of approximations are made in the calculation of the related visibility function, which is based on computing the averaged recombination history including only second order effects of baryon-density perturbations on the recombination dynamics [e.g., 18, 19, for more details of the model]. Indeed the non-linear nature of the ionization history responses to baryon density fluctuations suggests that this may be omitting additional effects, which in a more rigorous treatment could manifest in enhanced higher order moments of the Thomson visibility function, an aspect that may be worth exploring in the future.
@@ V Forecasts for SPT-3G
Figure 14: Power spectra residuals between predictions from the ModRec model and the Planck best-fit cosmology, taken as the reference cosmology. Also shown are binned data from Planck, as well as the forecasted bandpower errors for SPT-3G in the Main-1500 configuration. Each have been binned with $\Delta\ell=100$. The power spectra correspond to models having the labeled values of $H_{0}$, in the same manner as Fig. 8.
Figure 15: Fractional uncertainties in our recombination reconstructions for the Planck, Planck+BAO, and Planck+BAO+SPT data combinations. Dashed lines indicated the $68\%$ and $95\%$ confidence intervals. The color indicates the density of trajectories through a given point, as in Fig. 5. SPT data will decrease the our reconstruction uncertainties by roughly a factor of 2, relatively uniformly across the redshift range in question. The fiducial cosmology for the SPT mock likelihood was taken to be the Planck mean cosmology. Our prior forces deviations from the standard ionization history to vanish at $z_{\rm min}=533$ and $z_{\rm max}=1600$, leading to artificially tight constraints at these boundaries.
@@
$\log{\mathcal{L}}=\left(\mathbf{C}_{\ell}-\mathbf{C}_{\ell}^{\rm fid}\right)^{\intercal}\Sigma_{\ell}^{-1}\left(\mathbf{C}_{\ell}-\mathbf{C}_{\ell}^{\rm fid}\right)$ ) (16)
where $\mathbf{C}_{\ell}=\left(C_{\ell}^{TT},C_{\ell}^{EE},C_{\ell}^{TE},C_{\ell}^{\phi\phi}\right)$ and the Planck 2018 best-fit cosmology is taken as the fiducial. The covariance matrix $\Sigma_{\ell}$ incorporates the effects of both instrumental noise and foregrounds and is constructed in the manner described in [77]. We use noise levels corresponding to the Main-1500 configuration of SPT-3G, representing 5 years of observations on a sky area of 1500 $\rm deg^{2}$. This corresponds to the data that has already been collected since SPT-3G began observations in 2019.
As the observation regions for Planck and SPT-3G are not independent, we remove large scale modes from the mock SPT-3G data to avoid double counting information. The $\ell_{\rm min}$ threshold was determined as roughly the multipole at which the signal-to-noise for SPT-3G is twice that of Planck. This approach was validated in Prabhu et al. [77], and was found to give results less than 8% worse than a full Fisher accounting for the overlapping sky regions. The multipole ranges we used were $\ell\in[2000,3500],[1000,3500],[750,4000],[30,4000]$ for TT, TE, EE, and $\phi\phi$ respectively. We bin the data with $\Delta\ell=100$, and discard bins which are centered outside of these ranges. This mock data is used along with Planck CMB and BOSS/eBOSS BAO data in an MCMC.
@@
In Fig. 14, we show predictions for the same models presented as dashed lines in in Figs. 10 and  11. We first bin models from the Planck-only chain according to their value of $H_{0}$ and select only models which fit better than a $\Lambda$CDM best fit, plotting the mean from each bin. We see that even for $H_{0}\approx 68$ km/s/Mpc, the ModRec model power spectrum predictions differ from those of $\Lambda$CDM. This indicates that the new freedom introduced in recombination improves the fit relative to the case of standard recombination — in the ModRec model, the standard recombination history does not provide the best fit even for models near the Planck best fit assuming $\Lambda$CDM. This is also reflected by the fact that the mean ionization fraction for the Planck chain in Fig. 5 differs from the standard prediction. There are deviations from the $\Lambda$CDM best fit (constrained by Planck) at the level of a few percent in the damping tail for the TT and EE spectra. The largest deviation from the $\Lambda$CDM reference is seen in the cross-spectrum, where all of the predictions deviate from the Planck prediction.
Taking the $H_{0}=71$ km/s/Mpc bin mean as a representative example of a high-$H_{0}$ recombination history, we find that $\chi^{2}_{\rm SPT}=205$. The binned mock SPT data, after large scale modes are removed, consists of 113 data points which are constraining the 13 parameters of the ModRec model. The probability for $\chi^{2}_{\rm SPT}$ to exceed 205 for this mock data is approximately $2\times 10^{-9}$, indicating that forthcoming SPT data will be able to rule out (or confirm) such a deviation at extremely high confidence  [^3] For a Gaussian distribution, this would correspond to a $5.9\sigma$ detection, although we note that the ModRec posteriors are highly non-Gaussian.. Precise measurements of the temperature and polarization power spectra at small scales will therefore be important for constraining alternative models of the recombination era.